Circular No. 6414

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

PSR 1055–52 R. Mignani, P. A. Caraveo, and G. F. Bignami, Istituto di Fisica Cos-

mica, Milan, report observing the field of this radio/x-ray/γ-ray pulsar on May 11 with the Faint Object Camera onboard the Hubble Space Telescope: “The observation was performed in the F342 wideband filter centered at 340.2 nm (Δλ = 70.2 nm) for a total exposure time of 8900 s. We have identified a point source of mag 24.88 ± 0.1 at α = 10^{h}57^{m}58^{s}.83, δ = −52^{o}26′56′′.3 (equinox 2000.0), coincident within 0′′.1 with the radio-pulsar position. The object’s flux appears compatible with the low-energy extrapolation of the ROSAT soft-x-ray spectrum — i.e., a blackbody curve with a temperature of 750 000 K and a hydrogen column density of 6 × 10^{20}.”

GRO J1744–28 K. Jahoda, Goddard Space Flight Center (GSFC); T. Strohmayer and

R. Corbet, GSFC and Universities Space Research Association; and M. Stark, GSFC and University of Maryland, report: “GRO J1744–28 resumed bursting activity between May 29 and June 1. The Rossi X-ray Timing Explorer (RXTE) Proportional Counter Array (PCA) observed three bursts during a 151-min observation on June 1, and seven bursts during a 179-min observation on June 2. The average power spectrum of the June 2 bursts shows 6-percent rms variability at the 2.1-Hz pulse period of GRO J1744– 28, indicating that these bursts come from this source, and the peak power at 2.1 Hz reaches 10 percent rms. The brightest burst was 9000 counts/s (0.7 Crab). All of these bursts have rise times less than a few seconds; while most of the bursts have durations of 20–30 s, some continue up to 100 s. The burst spectrum can be characterized by a hard power-law (photon index 1.3) cutoff at ∼ 20 keV with a 10-keV e-folding energy, similar to the spectrum previously measured for GRO J1744–28. While a nearby source hampers measurements of the quiescent flux, the nonbursting pulsed flux is 13 ± 0.5 counts/s, an increase of at least a factor of two since May 29. This suggests that the quiescent flux has also increased. RXTE will continue to observe GRO J1744–28 regularly.”

COMET 22P/KOPFF Visual m_{1} estimates: May 18.55 UT, 8.8 (S. O’Meara, Volcano, HI,

10×35 binoc.); June 2.72, 8.8 (G. Wolf, Wellington, N.Z., 0.12-m refractor).

1996 June 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green